[The following information applies to the questions displayed below.]

BMX Company has one employee. FICA Social Security taxes are 6.2% of the first $127,200 paid to its employee, and FICA Medicare taxes are 1.45% of gross pay. For BMX, its FUTA taxes are 3.1% and SUTA taxes are 2.9% of the first $7,000 paid to its employee.

Gross Pay through August 31 Gross Pay for September
a. $ 6,400 $ 800
b. 2,000 2,100
c. 120,900 8,000

Compute BMX’s amounts for each of these four taxes as applied to the employee’s gross earnings for September under each of three separate situations (a), (b), and (c). (Round your answers to 2 decimal places. Leave no cells blank.)

a)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security
FICA-Medicare
FUTA
SUTA
b)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security
FICA-Medicare
FUTA
SUTA
c)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security
FICA-Medicare
FUTA
SUTA
Answer #1
a)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 800 6.20% 49.60
FICA-Medicare 800 1.45% 11.60
FUTA 600 3.10% 18.60
SUTA 600 2.90% 17.40
b)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 2100 6.20% 130.20
FICA-Medicare 2100 1.45% 30.45
FUTA 2100 3.10% 65.10
SUTA 2100 2.90% 60.90
c)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 6300 6.20% 390.60
FICA-Medicare 8000 1.45% 116.00
FUTA 0 3.10% 0.00
SUTA 0 2.90% 0.00
Note: Earnings taxable for FUTA and SUTA in (a) = 7000-6400 = $600
Note: Earnings taxable for Social Security in (c) = 127200-120900 = $6300